# Low-Dimensional Adaptation of Rectified Flow: A New Perspective through the Lens of Diffusion and Stochastic Localization

### v1

@@ Abstract
In recent years, Rectified flow (RF) has gained considerable popularity largely due to its generation efficiency and state-of-the-art performance. In this paper, we investigate the degree to which RF automatically adapts to the intrinsic low dimensionality of the support of the target distribution to accelerate sampling. We show that, using a carefully designed choice of the time-discretization scheme and with sufficiently accurate drift estimates, the RF sampler enjoys an iteration complexity of order $O(k/\varepsilon)$ (up to log factors), where $\varepsilon$ is the precision in total variation distance and $k$ is the intrinsic dimension of the target distribution. In addition, we show that the denoising diffusion probabilistic model (DDPM) procedure is equivalent to a stochastic version of RF by establishing a novel connection between these processes and stochastic localization. Building on this connection, we further design a stochastic RF sampler that also adapts to the low-dimensionality of the target distribution under milder requirements on the accuracy of the drift estimates, and also with a specific time schedule. We illustrate with simulations on the synthetic data and text-to-image data experiments the improved performance of the proposed samplers implementing the newly designed time-discretization schedules. Codes are available in https://github.com/roysaptaumich/Low-dim-RF
@@ 1 Introduction
In the past few years, Rectified flow or Flow matching (Liu, 2022; Lipman et al., 2022) has gained much attention due to its state-of-the-art generative performance across various modalities including image generation (Liu et al., 2024; Yan et al., 2024), audio generation (Wang et al., 2024; Liu et al., 2025a), and video generation (Chen et al., 2025; Liu et al., 2025b). The success of RF stems from two key factors: (1) RF is a deterministic ordinary differential equation (ODE) based sampler that is easier to simulate compared to other stochastic samplers, and (2) RF learns almost straight trajectories between source and target data distribution via multiple reflow steps which leads to faster sampling. However, RF still faces several challenges.
  • The reflow process is computationally intensive as it requires a large number of generated pairs to model the target distribution.
  • It is well known that the multiple reflow steps can quickly lead to performance degradation of the learned flow model (Zhu et al., 2024; Kim et al., 2025).
Therefore, understanding and mitigating the learning error in the first reflow itself is extremely important to improve generation quality. There has been prior works (Potaptchik et al., 2024; Azangulov et al., 2024; Huang et al., 2024; Li and Yan, 2024; Liang et al., 2025) that shows the low-dimensional adaptation of DDPM. Therefore, DDPM enjoys accelerated convergence to achieve improved sampling which is consistent with empirical findings even when the ambient dimension is very high. To the best of our knowledge such results are not currently present for RF models. A recent work by Lee et al. (2024) has empirically shown that a certain U-shaped time-discretization improves the generative quality of the first reflow model. However, theoretical understanding of such phenomenon remains elusive and necessitates in-depth investigation.
To this end, we make three key contributions in this paper:
@@
    We design a novel U-shaped time-discretization that provably allows any off-the-shelf pretrained RF sampler to quickly achieve improved generation quality by adapting to intrinsic low-dimensional structure of the target distribution. In particular, the RF sampler enjoys $O(k/\varepsilon)$ (upto log factors) iteration complexity, where $\varepsilon$ is the precision of the generated samples in total variation distance and $k$ is the intrinsic dimension of the target distribution.
@@
    We demonstrate that DDPM is equivalent to a stochastic version of RF (which we refer to as Stoc-RF) under a proper time change by establishing a novel connection between these processes and stochastic localization (Eldan, 2020), which has been recently shown to be provide a unifying and effective framework for representing and analyzing diffusion-based sampling processes (Benton et al., 2023; Montanari, 2023; El Alaoui and Montanari, 2022). This equivalence is convenient as it allows us to view Stoc-RF through the lens of well understood DDPM processes, which are known to adapt to the low-dimensional structure of the target distribution. In fact, our analysis shows that stochastic localization can be leveraged to provide a unified and convenient representation of a variety of sampling methods.
@@
    Using the connection between DDPM and Stoc-RF, we further design a stochastic RF sampler that provably also adapts to the low-dimensional structure of the target distributions under less stringent assumptions on the learned model. We illustrate the superior performance of this sampler with various simulations. This also addresses one of the open questions posed in the ICML tutorial talk Liu (2025).
Notation. Throughout the paper, with some abuse of notations, we will identify a probability distribution with its Lebesgue density, which we will assume exists unless otherwise noted. For a distribution (Lebesgue density) $p$, we write ${\sf supp}(p)$ to denote the support of $p$. For a random variable $X$ we use $\text{Law}(X)$ to denote its distribution. For two distributions $P$ and $Q$ on $\mathbb{R}^{d}$ with Lebesgue densities $p$ and $q$ respectively, their total variation distance is denoted by ${\sf TV}(p,q):=\sup_{A}|{P(A)-Q(A)}|=1/2\int_{\mathbb{R}^{d}}|p(x)-q(x)|dx$ , where the supremum is over all Borel subsets of $\mathbb{R}^{d}$.
@@ 2.1 Rectified Flow
In this section, we briefly introduce the basics of Rectified flow, a generative model that transitions between two distributions $p_{0}$ and $p_{1}$ by solving ODEs. We refer the reader to Liu (2024, 2022) for an exhaustive treatment. Let $p_{\rm{data}}:=p_{1}=\text{Law}(X_{1})$ be the target data distribution on $\mathbb{R}^{d}$, the data generating distributi from which we would like to draw samples. The linear-interpolation process is defined by by
@@
where the minimization is over all functions $f:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$. For each $t$, The above objective in (2) is minimized at
$(x,t)\mapsto v_{t}(x):=v(x,t)$ $=\mathbb{E}[X_{1}-X_{0}\mid X_{t}=x].$ (3)
Under appropriate regularity conditions (investigated in Bansal et al., 2025; Mena et al., 2025), the ODE
@@
$\frac{\partial p_{t}}{\partial t}+\nabla.(v_{t}p_{t})=0.$
Hence, ODE (4) can be used for sampling from the target distribution $p_{1}$.
@@ 2.2 Stochastic rectified flow (Stoc-RF)
Since the work of DDIM (Song et al., 2021a) and DDPM (Song et al., 2021b), it is well known that SDEs can be converted to ODEs to obtain deterministic samplers. Conversely, it is also possible to convert ODEs to its equivalent SDE form to obtain stochastic samplers. In fact, prior works (Albergo et al., 2023; Xue et al., 2024; Hu et al., 2025a) have shown that stochastic samplers typically enjoy better generation quality. One reason behind the inferior performance of deterministic samplers is the possibility of error accumulation during the sampling step along the discrete trajectories stemming from inaccurate velocity estimates. In other words, if $Y_{t_{i}}$ differs significantly from $Z_{t_{i}}$ for some $t_{i}$, then the Euler step (5) accumulates error due to inaccurate evaluation of $\widehat{v}_{t_{i}}(Y_{t_{i}})$. Therefore, this can further reinforce the deviation in evaluating $Y_{t_{i+1}}$. A way to mitigate the compounding error from successive Euler steps is to introduce a stochastic Langevin correction at each step. Empirically, (Hu et al., 2025b) showed that the resulting stochastic RF samplers enjoy better text rendering quality over deterministic samplers. We further illustrate this point with Figure 1, showing that the ODE-based RF trajectories appear to generate more outliers compared to the SDE-based Stoc-RF sampler.
@@
Following e.g. Liu (2024, Chapter 5), the stochastic version of RF is specified by SDE
@@
where $\tilde{v}_{t}(x)=v_{t}(x)+\gamma_{t}s_{t}(x)-\gamma_{t}\nabla\log\tilde{p}_{t}(x)$. Since the continuity equations (2.1) and (7) coincide, $\tilde{p}_{t}=p_{t}$ , i.e., $\text{Law}(\tilde{Z}_{t})=\text{Law}(Z_{t})$, for all $t$, demonstrating that the SDE (6) can indeed be used for sampling from $p_{1}$. In practice, this is accomplished via an appropriate time discretization scheme
@@
where $\left\{{t_{i}}\right\}_{i=0}^{N}$ is an appropriate schedule of times, the $\xi_{t_{i}}$’s are i.i.d. draws from $N(0,I_{d})$, and the $\tilde{\eta}_{i}$’s are (typically small) step-sizes. The Langevin term essentially acts as a course-corrector to the trajectory of (8) to adjust the trajectory distribution closer to $p_{t}$. Finally, one has freedom to choose the diffusion coefficients $\gamma_{t}$. In this paper we will restrict ourselves to the choice $\gamma_{t}=(1-t)/t$, which, as we will see below, is the appropriate noise scaling that yields equivalence between the Stoc-RF SDE (6) and DDPM through the lens of Stochastic localization. This also addresses an open question raised during the ICML tutorial talk Liu (2025) regarding the motivation of such choice.
3 Stochastic localization, and equivalence between RF and DDPM
@@
Stochastic localization is a probabilistic technique to investigate the properties of high-dimensional probability distributions due to (Eldan, 2020) that provides a powerful, unifying mathematical framework for understanding, analyzing, and improving score-based generative models and sampling algorithms. See El Alaoui and Montanari (2022), Montanari (2023) and the survey Shi et al. (2025), to which we refer the reader. To sample from the target data distribution $p_{1}$, SL focusses on an appropriate measured-valued stochastic process $(\nu_{s})_{s\geq 0}$ such that $\nu_{s}$ “localizes” towards $p_{1}$ as $s\to\infty$. A straightforward way to formulate this process is by letting, for any $s\geq 0$, $\nu_{s}:=\text{Law}(X_{1}\mid U_{s})$, where $X_{1}\sim p_{1}$ and
@@
with $(B_{s})_{s\geq 0}$ a standard Brownian motion in $\mathbb{R}^{d}$ (Montanari, 2023), independent of $X_{1}$. Thus, $\nu_{s}$ is a random measure depending on $U_{s}$. Note that $U_{s}/s\to X_{1}$ almost surely as $s\to\infty$, i.e., $\nu_{s}$ does indeed localize around $p_{1}$.
While enlightening, the construction (9) does not yield a sampler from $p_{1}$, as it uses $X_{1}\sim p_{1}$ in the definition of $U_{s}$. Towards that goal, it is instead convenient to consider the alternative process $(\tilde{U}_{s})_{s\geq 0}$ defined as the solution of the SDE
@@
Lemma 3.1 (Eldan (2020), Equation (11)).
For all $s\geq 0$, $\frac{d}{ds}\mathbb{E}(\mathbf{A}_{s})=-\mathbb{E}(\mathbf{A}_{s}^{2})$.
This is one of the crucial results that we will use later to control the discretization error of ODE and SDE samplers of RF.
@@
where $\beta(\tau)$ is the diffusion coefficient and $\left\{{B^{\prime}_{\tau}}\right\}_{\tau\geq 0}$ is standard Brownian motion. If $\beta(\tau)=1$, then the SDE (11) reduces to the classical Ornstein-Ulhenbeck (OU) process. Typically, The forwards process is continued till a large time $\tau=N$ such that $Y_{N}^{\prime}$ is approximately Gaussian. Let $q_{\tau}(y_{\tau}^{\prime})$ be the marginals of $y_{\tau}^{\prime}$ following (11). Then in order to sample from the target distribution $p_{1}$, a reverse process $(\overleftarrow{Y}_{\tau})_{\tau\in[0,N]}$ is formulated, starting at $\overleftarrow{Y}_{0}=Y_{N}^{\prime}$ and evolving according to the reverse-time SDE
$d\overleftarrow{Y}_{\tau}=$ $\{\overleftarrow{Y}_{\tau}+2\nabla\log q_{N-\tau}(\overleftarrow{y}_{\tau})\}\beta(N-\tau)d\tau+\sqrt{2\beta(N-\tau)}d\overleftarrow{B}_{\tau},$ (12)
where $(\overleftarrow{B}_{\tau})_{\tau\geq 0}$ is a standard Brownian motion. Classical results in SDE (Anderson, 1982; Haussmann and Pardoux, 1986) show that $\overleftarrow{Y}_{\tau}\overset{d}{=}Y_{N-\tau}^{\prime}$ which allows us to generate samples from $p_{1}$ via simulating the backward SDE (12).
@@
where $(W_{\tau})_{\tau\geq 0}$ is a Brownian motion. Note that $\text{Law}(\tilde{X}_{t})=\text{Law}(X_{t})$, for all $t\in[0,1]$. In fact, as the next result show, the entire process (13) has the same law of the SL process (9), up to time-change.
@@
The proof is deferred to Appendix C.1. In light of Lemma 3.2 and Lemma 3.3, we immediately conclude that
@@
Lemma 3.3 and the equivalence result in (14) can be extended to cover general stochastic interpolation (Albergo and Vanden-Eijnden, 2023) schemes beyond RF. In detail, instead of a linear interpolation, one can consider a general deterministic affine interpolation process $I_{\theta}=a_{\theta}X_{1}+b_{\theta}X_{0}$, where $a_{(\cdot)}$ and $b_{(\cdot)}$ are smooth functions of $\theta\in[0,1]$, with $a_{0}=b_{1}=0$, and $a_{1}=b_{0}=1$. When $X_{0}\sim N(0,I_{d})$, the associated stochastic interpolant is the process $(\tilde{I}_{\theta})_{\theta\in[0,1]}$, where $\tilde{I}_{\theta}=a_{\theta}X_{1}+a_{\theta}W_{r^{2}_{\theta}}$, with $r_{\theta}=b_{\theta}/a_{\theta}$ and $\left\{{W_{t}}\right\}_{t\geq 0}$ a Brownian motion. It is easy to see that $\tilde{I}_{\theta}\overset{d}{=}I_{\theta}$, for each $\theta$. Moreover, each stochastic interpolant can be rescaled and time-transformed to return the SL process, in a sense made precise by the next result.
@@
In particular, in virtue of Lemma 3.3, RF is equal in law to any interpolant process $I_{\theta}$ under the time change $t(\theta)=\frac{a_{\theta}}{a_{\theta}+b_{\theta}}$. This equivalence among stochastic interpolants is not new, as it has also been shown in Liu (2024) by direct arguments. Connection between diffusion and flow matching has also been studied in Albergo et al. (2023); Gao et al. (2024); Ma et al. (2024) through different perspectives. By leveraging the representation properties of the SL process, we establish this result in a more unified manner, which has the added benefit of directly applying to both stochastic interpolants and SDEs. The proof of Proposition 3.5 is deferred to Appendix C.3.

### v2

# Low-Dimensional Adaptation of Rectified Flow: A Diffusion and Stochastic Localization Perspective

@@ Abstract
In recent years, Rectified flow (RF) has gained considerable popularity largely due to its generation efficiency and state-of-the-art performance. In this paper, we investigate how well RF automatically adapts to the intrinsic low dimensionality of the support of the target distribution to accelerate sampling. We show that, using a carefully designed choice of the time-discretization scheme and with sufficiently accurate drift estimates, the RF sampler enjoys an iteration complexity of order $O(k/\varepsilon)$ (up to log factors), where $\varepsilon$ is the precision in total variation distance and $k$ is the intrinsic dimension of the target distribution. In addition, we show that the denoising diffusion probabilistic model (DDPM) procedure is equivalent to a stochastic version of RF by establishing a novel connection between these processes and stochastic localization. Building on this connection, we further design a stochastic RF sampler that also adapts to the low-dimensionality of the target distribution under mild requirements on the accuracy of the drift estimates, and also with a specific time schedule. We illustrate the efficacy of newly designed time-discretization schedules with simulations on the synthetic data and text-to-image (T2I) data experiments.
@@ 1 Introduction
Recently, Rectified flow or Flow matching (Liu, 2022; Lipman et al., 2022) has gained much attention due to its state-of-the-art generative performance across various modalities including image (Liu et al., 2024; Yan et al., 2024), audio (Wang et al., 2024; Liu et al., 2025a), and video (Chen et al., 2025; Liu et al., 2025b). Its effectiveness arises from a deterministic ordinary differential equation (ODE) based sampler and the ability of reflow to learn nearly straight transport paths, resulting in faster sampling. This has naturally sparked theoretical interests (Bansal et al., 2025; Guan et al., 2025; Zhou and Liu, 2025) to understand the generation quality of RF from a convergence viewpoint under a suitable metric. However, these recent results suffer from curse of dimensionality. Therefore, they fail to explain the impressive performance of RF for high-dimensional data, e.g., image data, even though they reside in a low-dimensional space (Ansuini et al., 2019; Pope et al., 2021). In contrast, recent works (Potaptchik et al., 2024; Azangulov et al., 2024; Huang et al., 2024; Li and Yan, 2024; Liang et al., 2025) have established low-dimensional adaptation of DDPM and denoising diffusion implicit models (DDIM) during sampling. This explains the accelerated and high-quality generative capabilities of DDPM and DDIM in practice, even for high-dimensional data. To the best of our knowledge such results are not currently present for RF models. To this end, we make three key contributions in this paper:
@@
    We design a novel U-shaped time-discretization that provably allows RF to quickly achieve improved generation quality in the first reflow process by adapting to the intrinsic low-dimensionality of the target distribution. This aligns with the recent work by Lee et al. (2024), which has empirically shown that a certain U-shaped time-discretization improves the generative quality of the first reflow model. In particular, the RF sampler enjoys $O(k/\varepsilon)$ (upto log factors) iteration complexity, where $\varepsilon$ is the precision of the generated samples in total variation distance and $k$ is the intrinsic dimension of the target distribution. This is particularly useful because it can eliminate repeated reflow steps, which are computationally costly and often cause model collapse (Zhu et al., 2024; Kim et al., 2025).
@@
    Using a novel stochastic localization (Eldan, 2013, 2020) perspective, we demonstrate that DDPM is equivalent to the stochastic version of RF (which we refer to as STOC-RF) under a proper time change. Stochastic localization has been recently shown to be provide a unifying and effective framework for representing and analyzing diffusion-based sampling processes (Benton et al., 2023; Montanari, 2023; El Alaoui and Montanari, 2022). This equivalence is convenient as it allows us to view STOC-RF through the lens of well-understood DDPM processes, which are known to adapt to the low-dimensionality of the target distribution. In fact, our analysis shows that stochastic localization can be leveraged to provide a unified and convenient representation of a variety of sampling methods.
@@
    Using the connection between DDPM and STOC-RF, we further design a stochastic RF sampler that provably also adapts to the low-dimensional structure of the target distributions under less stringent assumptions on the learned model. We illustrate the superior performance of this sampler with various simulations. This also addresses one of the themes discussed in the ICML tutorial talk by Liu (2025).
Notation. Throughout the paper, with some abuse of notations, we will identify a probability distribution with its Lebesgue density, which we will assume exists unless otherwise noted. For two distributions $P$ and $Q$ on $\mathbb{R}^{d}$ with Lebesgue densities $p$ and $q$ respectively, their total variation distance is denoted by ${\sf TV}(p,q):=\sup_{A}|{P(A)-Q(A)}|=1/2\int_{\mathbb{R}^{d}}|p(x)-q(x)|dx$ , where the supremum is over all Borel subsets of $\mathbb{R}^{d}$.
@@ 2.1 Rectified Flow
In this section, we briefly introduce the basics of Rectified flow, a generative model that transitions between two distributions $p_{0}$ and $p_{1}$ by solving ODEs. We refer the reader to Liu (2024, 2022) for an exhaustive treatment. We let $p_{1}:=\text{Law}(X_{1})$ be the target data distribution with its support $\mathcal{X}\subseteq\mathbb{R}^{d}$, the data generating distribution from which we would like to draw samples. The linear-interpolation process is defined by by
@@
where the minimization is over all functions $f:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$. For each $t$, The above objective in (2) is minimized at $(x,t)\mapsto v_{t}(x):=v(x,t)=\mathbb{E}[X_{1}-X_{0}\mid X_{t}=x].$ Under appropriate regularity conditions (investigated in Bansal et al., 2025; Mena et al., 2025), the ODE
@@
$\frac{\partial p_{t}}{\partial t}+\nabla.(v_{t}p_{t})=0.$ (4)
Hence, ODE (3) can be used for sampling.
@@ 2.2 Stochastic rectified flow (STOC-RF)
Since the work of DDIM (Song et al., 2021a) and DDPM (Song et al., 2021b), it is well known that SDEs can be converted to ODEs to obtain deterministic samplers. Conversely, it is also possible to convert ODEs to its equivalent SDE form to obtain stochastic samplers. In fact, prior works (Albergo et al., 2023; Xue et al., 2024; Hu et al., 2025a) have shown that stochastic samplers typically enjoy better generation quality. One reason behind the inferior performance of deterministic samplers is the possibility of error accumulation during the sampling step along the discrete trajectories stemming from inaccurate velocity estimates. In other words, if $Y_{t_{i}}$ differs significantly from $Z_{t_{i}}$ for some $t_{i}$, then the Euler step (5) accumulates error due to inaccurate evaluation of $\widehat{v}_{t_{i}}(Y_{t_{i}})$. Therefore, this can further reinforce the deviation in evaluating $Y_{t_{i+1}}$. A way to mitigate the compounding error from successive Euler steps is to introduce a stochastic Langevin correction at each step. Empirically, Hu et al. (2025b) showed that the resulting stochastic RF samplers enjoy better text rendering quality over deterministic samplers. We further illustrate this point with Figure 1, showing that the ODE-based RF trajectories appear to generate more outliers compared to the SDE-based STOC-RF sampler.
@@
Following e.g. Liu (2024, Chapter 5), the stochastic version of RF is specified by the SDE
@@
where $\tilde{v}_{t}(x)=v_{t}(x)+\gamma_{t}s_{t}(x)-\gamma_{t}\nabla\log\tilde{p}_{t}(x)$. Since the continuity equations (4) and (7) coincide, $\tilde{p}_{t}=p_{t}$ , i.e., $\text{Law}(\tilde{Z}_{t})=\text{Law}(Z_{t})$, for all $t$, demonstrating that the SDE (6) can indeed be used for sampling from $p_{1}$. In practice, this is accomplished via an appropriate time discretization scheme
@@
where $\left\{{t_{i}}\right\}_{i=0}^{N}$ is an appropriate schedule of times, the $\xi_{t_{i}}$’s are i.i.d. draws from $N(0,I_{d})$, and the $\tilde{\eta}_{i}$’s are (typically small) step-sizes. The Langevin term essentially acts as a course-corrector to the trajectory of (8) to adjust the trajectory distribution closer to $p_{t}$. Finally, one has freedom to choose the diffusion coefficients $\gamma_{t}$. In this paper we will restrict ourselves to the choice $\gamma_{t}=(1-t)/t$, which, as we will see below, is the appropriate noise scaling that yields equivalence between the STOC-RF SDE (6) and DDPM through the lens of Stochastic localization. This also addresses an question raised during the ICML tutorial talk Liu (2025) regarding the motivation of such choice.
3 Stochastic localization, RF and DDPM
@@
SL is a probabilistic technique to study the properties of high-dimensional distributions due to Eldan (2013, 2020) that provides a powerful, unifying framework for understanding, analyzing, and improving score-based generative models and sampling algorithms. See El Alaoui and Montanari (2022), Montanari (2023) and the survey by Shi et al. (2025). To sample from the target data distribution $p_{1}$, SL focuses on an appropriate measured-valued stochastic process $\{\nu_{s}\}_{s\geq 0}$ such that $\nu_{s}$ “localizes” towards $p_{1}$ as $s\to\infty$. A straightforward way to formulate this process is by letting, for any $s\geq 0$, $\nu_{s}:=\text{Law}(X_{1}\mid U_{s})$, where $X_{1}\sim p_{1}$ and
@@
with $\{B_{s}\}_{s\geq 0}$ a standard Brownian motion in $\mathbb{R}^{d}$ (Montanari, 2023), independent of $X_{1}$. Thus, $\nu_{s}$ is a random measure depending on $U_{s}$. Note that $U_{s}/s\to X_{1}$ almost surely as $s\to\infty$, i.e., $\nu_{s}$ does indeed localize around $p_{1}$.
While enlightening, the construction (9) does not yield a sampler from $p_{1}$, as it uses $X_{1}\sim p_{1}$ in the definition of $U_{s}$. Therefore, it is instead convenient to consider the alternative process $\{\tilde{U}_{s}\}_{s\geq 0}$ defined as the solution of the SDE
@@
Lemma 3.1.
For all $s\geq 0$, $\frac{d}{ds}\mathbb{E}\left[{\mathbf{A}_{s}}\right]=-\mathbb{E}\left[{\mathbf{A}_{s}^{2}}\right]$.
This is a crucial result for controlling the discretization error of ODE and SDE samplers of RF.
@@
where $\beta(\tau)$ is the diffusion coefficient and $\left\{{B^{\prime}_{\tau}}\right\}_{\tau\geq 0}$ is standard Brownian motion. If $\beta(\tau)=1$, then the SDE (11) reduces to the classical Ornstein-Ulhenbeck (OU) process. Typically, The forwards process is continued till a large time $\tau=N$ such that $Y_{N}^{\prime}$ is approximately Gaussian. Let $q_{\tau}(y_{\tau}^{\prime})$ be the marginals of $y_{\tau}^{\prime}$ following (11). Then in order to sample from the target distribution $p_{1}$, a reverse process $\{\overleftarrow{Y}_{\tau}\}_{\tau\in[0,N]}$ is formulated, starting at $\overleftarrow{Y}_{0}=Y_{N}^{\prime}$ and evolving according to the reverse-time SDE
$d\overleftarrow{Y}_{\tau}=$ $\{\overleftarrow{Y}_{\tau}+2\nabla\log q_{N-\tau}(\overleftarrow{y}_{\tau})\}\beta(N-\tau)d\tau$ (12)
$+\sqrt{2\beta(N-\tau)}d\overleftarrow{B}_{\tau},$
where $\{\overleftarrow{B}_{\tau}\}_{\tau\geq 0}$ is a standard Brownian motion. Classical results in SDE (Anderson, 1982; Haussmann and Pardoux, 1986) show that $\overleftarrow{Y}_{\tau}\overset{d}{=}Y_{N-\tau}^{\prime}$ which allows us to generate samples from $p_{1}$ via simulating the backward SDE (12).
@@
where $\{W_{\tau}\}_{\tau\geq 0}$ is a Brownian motion. Note that $\text{Law}(\tilde{X}_{t})=\text{Law}(X_{t})$, for all $t\in[0,1]$. In fact, as the next result show, the entire process (13) has the same law of the SL process (9), up to time-change.
@@
The proof is in Appendix C.1. Crucially, Lemma 3.3 yields the following Corollary to Lemma 3.1 which is used to control the discretization error of the RF sampler.
Corollary 3.4.
Let ${\boldsymbol{\Sigma}}_{t}:=\text{Cov}(X_{1}\mid X_{t})$ for $t\in[0,1)$. Then $\frac{d}{dt}\mathbb{E}({\boldsymbol{\Sigma}}_{t})=-\frac{2t}{(1-t)^{3}}\mathbb{E}({\boldsymbol{\Sigma}}_{t}^{2})$.
See Section D.1 for complete proof. Additionally, Lemma 3.2 and Lemma 3.3 leads to the following:
@@
Lemma 3.3 and the equivalence result in (14) can be extended to cover general stochastic interpolation (Albergo and Vanden-Eijnden, 2023; Albergo et al., 2023) schemes beyond RF. In detail, instead of a linear interpolation, one can consider a general deterministic affine interpolation process $I_{\theta}=a_{\theta}X_{1}+b_{\theta}X_{0}$, where $a_{(\cdot)}$ and $b_{(\cdot)}$ are smooth functions of $\theta\in[0,1]$, with $a_{0}=b_{1}=0$, and $a_{1}=b_{0}=1$. When $X_{0}\sim N(0,I_{d})$, the associated stochastic interpolant is the process $(\tilde{I}_{\theta})_{\theta\in[0,1]}$, where $\tilde{I}_{\theta}=a_{\theta}X_{1}+a_{\theta}W_{r^{2}_{\theta}}$, with $r_{\theta}=b_{\theta}/a_{\theta}$ and $\left\{{W_{t}}\right\}_{t\geq 0}$ a Brownian motion. It is easy to see that $\tilde{I}_{\theta}\overset{d}{=}I_{\theta}$, for each $\theta$. Moreover, each stochastic interpolant can be rescaled and time-transformed to return the SL process, in a sense made precise by the next result.
@@
In particular, in virtue of Lemma 3.3, RF is equal in law to any interpolant process $I_{\theta}$ under the time change $t(\theta)=\frac{a_{\theta}}{a_{\theta}+b_{\theta}}$. This equivalence among stochastic interpolants is not new, as it has also been shown in Liu (2024) by direct arguments. Connection between diffusion and flow matching has also been studied in Albergo et al. (2023); Albergo and Vanden-Eijnden (2022); Gao et al. (2024); Ma et al. (2024); Rout et al. (2025) through different perspectives. By leveraging the representation properties of the SL process, we establish this result in a more unified manner, which has the added benefit of directly applying to both stochastic interpolants and SDEs. The proof of Proposition 3.6 is deferred to Appendix C.3.